OpenStudy (anonymous):

Fan & Medal! Please help me What is the axis symmetry of Y=-x^2+x-7?

OpenStudy (anonymous):

Use the formula x= -b/2a.

OpenStudy (anonymous):

The problem is set up like this.... Y=-x^2+x-7 Y=ax^2 +bx +c

OpenStudy (ybarrap):

When you use the quadratic formula, there is a term, \(x=\cfrac{-b}{2a}\pm~something\). This give you a hint because the plus or minus part tells you that \(x=\cfrac{-b}{2a}\) is right in the middle; hence, is the axis of symmetry. Note a=-1, b=1 and c=-7 in your equation. http://en.wikipedia.org/wiki/Quadratic_formula